# Vivid-VR: Distilling Concepts from Text-to-Video Diffusion Transformer for Photorealistic Video Restoration

### v2

@@ Abstract
Figure 1: Video restoration results on both real-world and AIGC videos. To mitigate drift during fine-tuning of the controllable generation pipeline, we propose a concept distillation training strategy that preserves both texture realism and temporal coherence in the generated videos. Leveraging this strategy, Vivid-VR achieves impressive texture realism and visual vividness. (Zoom-in for best view)
@@ 1 Introduction
Video restoration aims to recover lost textures, fine details, and structural information from low-quality input videos to produce high-quality ones. Traditional reconstruction-based methods typically employ CNNs [33, 19, 4, 5, 6] and Transformers [13, 14] to extract visual cues for quality enhancement. However, these approaches face inherent limitations due to insufficient prior knowledge and the ill-posed nature of the inverse problem, reconstructing high-quality textures directly from severely degraded inputs remains extremely challenging. While GAN-based methods [35, 34] can generate some textures to a certain extent, their generative capacity remains limited.
@@
To overcome this challenge, we propose a concept distillation training strategy that leverages synthetic data generated by a pre-trained T2V model. The proposed approach begins with a source video and its corresponding text description obtained through a vision-language model (VLM). We first corrupt the source video with noise, then employ the pre-trained T2V model to perform denoising while incorporating the text description. This process yields a video that encapsulates the T2V model’s semantic understanding of the textual concepts, ensuring perfect modal alignment between the generated video and text description in the T2V model’s latent space. By blending these synthesized data with real training samples during fine-tuning, our method successfully transfers the T2V model’s conceptual knowledge to the video restoration model, thereby mitigating the distribution drift problem while preserving both texture realism and temporal coherence.
@@ 2 Related Work
Reconstruction-based Video Restoration. Early video restoration approaches [3, 33, 20, 19, 4, 5, 6, 13, 14] focused mainly on architectural improvements, data simulation, and loss functions to reconstruct high-quality videos by extracting visual information directly from degraded inputs. CNN-based solutions have focused on employing various strategies for temporal information integration, including optical flow estimation [3, 20], deformable convolutions [33], bidirectional feature propagation [4], and optical flow-guided deformable alignment modules [5]. The emergence of Transformers in computer vision [28, 9] introduced new possibilities through attention mechanisms capable of modeling long-term spatio-temporal relationships, yielding significant performance improvements [13, 14]. However, these methods often rely on simplified degradation models, limiting real-world applicability. Subsequent works [34, 43] have introduced more complex degradation simulations to improve generalization. Another critical limitation stems from traditional L1/L2 loss functions, which tend to produce perceptually flat results. GAN-based frameworks [35, 34] are consequently adopted that incorporate adversarial training to preserve richer textural details. Despite these advances, methods relying solely on input-derived cues without strong priors still produce overly smoothed results when handling severely degraded content.
@@
Diffusion-based Video Restoration. In recent years, diffusion-based generative models [24, 22, 1, 40] have made significant progress, which introduce a new paradigm for restoration tasks. Initial explorations focused on image restoration [32, 42, 7]. Based on Stable Diffusion [24], StableSR [32] developed a trainable time-aware encoder to embed input image information into the generation process, transforming it into an image restoration model. SUPIR [42] achieved remarkable results through model scaling for image restoration. However, these approaches fundamentally lack temporal modeling capabilities, resulting in severe frame inconsistencies when directly applied to video sequences. Early solutions [47, 39] attempted to mitigate this through temporal enhancement techniques, such as incorporating trainable temporal layers into diffusion denoiser and VAE decoder [47], or implementing optical flow-based motion compensation [39], yet these post-hoc adjustments proved inadequate for ensuring robust temporal coherence. Diffusion Transformers (DiT) [21] marked a significant breakthrough, enabling T2V models [40] to generate high-quality videos with superior temporal stability. This progress inspired video restoration methods based on the T2V models. SeedVR [31] combines the shift-window attention mechanism with DiT to improve computational efficiency. STAR [37] designs a dynamic frequency loss function to improve fidelity. Concurrently, efforts to improve inference efficiency have led to the design of one-step diffusion models [30, 8]. Nevertheless, existing methods still exhibit noticeable gaps in texture realism and temporal consistency compared to native T2V models, due to distribution drift from imperfect multimodal alignment of training data. To bridge this gap, our work introduces a concept distillation training strategy that effectively preserves the texture and temproal quality of the base T2V model.
@@ 3.1 Model Architecture
ControlNet Pipeline. Given the text tokens, the visual tokens and the timestep embedding, DiT and ControlNet both perform $T$ denoising steps. DiT comprises $N$ DiT blocks, while ControlNet contains $N/6$ blocks initialized from DiT’s first $N/6$ ones. During denoising process, ControlNet’s visual tokens are integrated into DiT through $N$ proposed Dual-branch Connectors. For the $i^{th}$ connector, the fusion process of the control visual tokens is:
$\hat{x}^{i}=x^{i}+MLP(c^{\lfloor i/6\rfloor})+CA(x^{i},c^{\lfloor i/6\rfloor}),$ (1)
where $x^{i}$ denotes the visual tokens from the $i^{th}$ DiT block; $c^{\lfloor i/6\rfloor}$ represents the corresponding ControlNet block visual tokens aligned with the $i^{th}$ DiT block; $MLP(\cdot)$ and $CA(\cdot)$ are the MLP layer and cross attention module respectively; $\hat{x}^{i}$ is the fused visual tokens. After $T$ denoising steps, the visual tokens are unpatchified and fed into the VAE decoder to generate the final high-quality video output.
@@ 3.2 Concept Distillation Training Strategy
Training Data Collection. Effective training of DiT-based video restoration models demands extensive high-quality text-video pairs, which existing public datasets [26, 18, 38, 25] lack in both scale and diversity. To address this, we collected a large-scale video pool consisting of $3$ million videos with resolutions higher than $1024\times 1024$, frame rates higher than $24$, and durations higher than $2$ seconds. These videos cover a wide range of scenes, including portraits, natural landscapes, plants and animals, urban landscapes, etc. To ensure video quality, we further screened these videos using the no-reference video quality assessment metrics [36, 46] to remove low-quality videos. For the remaining high-quality videos, we generated text descriptions using CogVLM2-Video [40], maintaining consistency with CogVideoX1.5-5B’s configuration. The final curated multimodal training dataset comprises $500K$ text-video pairs with exceptional quality and variety.
@@
Concept Distillation. Due to limitations of the VLM captioner, the constructed text-video data pairs are not perfectly aligned (see Figure 3 top row). This may lead to distribution drift during fine-tuning, which degrades the texture and temporal quality of the generated videos. While developing a more accurate VLM captioner would solve this, it is costly. Our objective is to maintain the generative distribution of the T2V base model, which can be achieved by aligning video and text in the latent space of the T2V model. To this end, we employ the T2V model itself (i.e., CogVideoX1.5-5B) to perform video-to-video translation, where the process is guided by the text description. Specifically, given a text-video pair, we perturb the source video by adding noise with a standard deviation corresponding to $T/2$ time steps. We then apply CogVideoX1.5-5B to denoise the video over $T/2$ steps, conditioned on the text description, yielding a synthesized video that integrates the textual concepts. As illustrated in Figure 3 (second row), the generated video largely retains the source content, but some concepts have been modified to better align with those in the text description. We randomly extract text-video pairs from the constructed multimodal training dataset and employ the aforementioned process to generate $100K$ sample pairs. These generated samples are then combined with the original training dataset to facilitate fine-tuning of the control module in our DiT-based video restoration model.
@@
where $x_{0}$ is the high-quality video sampled from training dataset; $x^{text}$ and $x^{lq}$ are the corresponding text description and the synthesized low-quality video using the degradation model from [34]; $t$ and $\epsilon$ are the time step and noise; $x_{t}=\sqrt{\bar{\alpha_{t}}}x_{0}+\sqrt{1-\bar{\alpha_{t}}}\epsilon$ is the noised latent of $x_{0}$, and $\bar{\alpha_{t}}$ is the cumulative multiplication of the variance corresponding to time step $t$; $v_{\theta}$ denotes the networks of DiT and ControlNet (including the control feature projector and the connectors); $v$ is the optimization target, which is defined as $v=\sqrt{\bar{\alpha_{t}}}\epsilon-\sqrt{1-\bar{\alpha_{t}}}x_{0}$. During training, only the control feature projector, ControlNet, and connectors are trained, and other parameters remain frozen.
@@ 4.1 Implementation Details
Figure 5: Visual comparison results on temporal consistency. (a) displays the #5\#5 frame of the input video, while (b)-(g) present the outputs at frames 0, 55, and $10$, where “CD” denotes the proposed concept distillation. Vivid-VR demonstrates superior temporal coherence, as evident from the consistent structure of windows and doors throughout the sequence. (Zoom in on the red arrow area in each frame)
Table 2: Ablation studies of the proposed method on the UGC50 testset, where “FT” denotes “fine-tuning”, “CA” denotes “cross attention”, and (f) is the setting of the proposed Vivid-VR. The best performances is marked in bold text.
@@ 4.2 Quantitative Results
To evaluate the performance of the proposed algorithm, we compare Vivid-VR against state-of-the-art approaches, including reconstruction-based methods (Real-ESRGAN [34]), generative image restoration methods (SUPIR [42]), and generative video restoration methods (UAV [47], MGLD [39], STAR [37], DOVE [8], SeedVR-7B [31], SeedVR2-7B [30]). The evaluation covers synthetic (SPMCS [27], UDM10 [41], YouHQ40 [47]) and real-world (VideoLQ [6]) benchmarks. Furthermore, we constructe two testsets, containing real-world UGC videos (UGC50) and AIGC videos (AIGC50). For real-world and AIGC videos lacking of ground truth, we employed no-reference image (NIQE [17], MUSIQ [12], CLIP-IQA [29]) and video quality assessments (DOVER [36], MD-VQA [46]). For synthetic benchmarks, we supplemented these no-reference metrics with full-reference evaluations using PSNR, SSIM, and LPIPS [45].
Table 1 presents quantitative comparisons on $6$ benchmark testsets. The proposed Vivid-VR significantly outperforms existing methods in no-reference metrics, achieving the best results in almost all metrics. At the same time, we also note its advantages in full-reference metrics appear less pronounced. We argue that this arises primarily from the inherent limitations of these metrics, which often fail to align with human perceptual preferences. For example, the LPIPS values of Figure 4(g1) and (i1) are $0.3112$ and $0.4297$, respectively, while Figure 4(i1) is more prefered by human. This phenomenon becomes particularly evident in generative restoration scenarios where severe input degradation allows multiple plausible high-quality outputs, making full-reference metrics inadequate for quality assessment. This has also been mentioned in [42] and has been noticed by quality assessment studies [2, 11, 10].
@@ 4.3 Qualitative Results
In addition, Figure 5 shows visual comparison results on temporal consistency. As shown in Figure 5(g), the proposed Vivid-VR demonstrates superior coherence. For example, the structures of the windows and doors are well-consistent throughout the sequence. In contrast, SUPIR [42] shows frame-wise inconsistency as it is an image-based restoration approach (see Figure 5(c)). While STAR [37] and SeedVR-7B [31] leverage T2V frameworks, their fine-tuning-induced distribution drift compromises temporal consistency (see Figure 5(d)-(e)). Notably, Vivid-VR exhibits similar degradation without using the proposed concept distillation strategy (see Figure 5(f)).
@@ 5 Analysis and Discussions
We have shown that the proposed Vivid-VR performs favorably against state-of-the-art methods. To better understand the proposed algorithm, we perform further analysis on the key components.
@@ 5.1 Effect of the Control Feature Projector
The T2V base model is trained on high-quality video data, making direct use of latent features from low-quality videos detrimental to generation quality. To address this issue, we propose a control feature projector to remove degradation artifacts from the low-quality video latents. To validate its effectiveness, we conducted an ablation study by disabling this module while keeping other settings unchanged. As shown in Table 2 ((a) vs (f)), the proposed control feature projector is able to improve video restoration. SUPIR [42] attempts to tackle this challenge by independently fine-tuning the VAE encoder. However, this decoupled optimization creates feature incompatibility with subsequent DiT and ControlNet, leading to suboptimal results (Table 2 (b)). While this problem can be solved through joint VAE encoder and video restoration optimization [7], this approach is expensive to train. In contrast, our proposed lightweight control feature projector can achieve similar effects at lower computational cost.
Figure 6: Effectiveness of the dual-branch ControlNet connector, where “CA” stands for cross attention. (Zoom-in for best view)
Figure 7: Effectiveness of the concept distillation strategy, where “CD” stands for concept distillation. (Zoom-in for best view)
@@ 5.2 Effect of the Dual-Branch Connector
For the ControlNet connector, we propose a dual-branch architecture combining an MLP for feature mapping with a cross attention mechanism for dynamic feature retrieval. One may wonder whether this design helps video restoration. To answer this question, we conduct three ablation studies: 1) disabling the cross attention branch and leaving only an MLP which is a naive connector design; 2) disabling the MLP branch; 3) adopting the ZeroSFT connector [42]. Table 2 and Figure 6 show the comparative results of the ablation experiments. When the cross attention branch is disabled, the MLP connector does not perform well and produces results lacking in detail (see Table 2(c) and Figure 6(e)). When the MLP branch is disabled, the video restoration model fails to converge due to its exclusive selection of DiT-like features from control inputs. As a result, ControlNet cannot achieve effective control, resulting in output results that do not match the input content (see Figure 6(f)). These experiments demonstrate the necessity of our dual-branch design in ContorlNet connectors.
In addition, the results in Table 2(d) show that the performance of ZeroSFT connector [42] is inferior to our proposed dual-branch connector. Furthermore, the normalization operation in ZeroSFT architecture often causes residual artifacts of adjacent frames to appear in the output frames (see Figure 6(d)), while removing the normalization leads to gradient explosion during training. In contrast, our proposed connector avoids these problems.

### v3

@@ Abstract
Figure 1: Video restoration results on both real-world and AIGC videos. To mitigate drift during fine-tuning of the controllable generation pipeline, we propose a concept distillation strategy that preserves both texture realism and temporal coherence in generated videos. Leveraging this strategy, Vivid-VR achieves impressive texture realism and visual vividness. (Zoom-in for best view)
@@ 1 Introduction
Video restoration aims to recover lost textures, fine details, and structural information from low-quality (LQ) input videos to produce high-quality (HQ) ones. Traditional reconstruction-based methods typically employ CNNs [32, 19, 4, 5, 6] and Transformers [13, 14] to extract visual cues for quality enhancement. However, these approaches face inherent limitations due to insufficient prior knowledge and the ill-posed nature of the inverse problem, reconstructing high-quality textures directly from severely degraded inputs remains extremely challenging. While GAN-based methods [34, 33] can generate some textures to a certain extent, their generative capacity remains limited.
@@
To overcome this challenge, we propose a concept distillation training strategy that leverages synthetic data generated by a pre-trained T2V model. The proposed approach begins with a source video and its corresponding text description obtained through a vision-language model (VLM). We first corrupt the source video with noise, then employ the pre-trained T2V model to perform denoising while incorporating the text description. This process yields a video that encapsulates the T2V model’s semantic understanding of the textual concepts, ensuring inherent modal alignment between the generated video and text description in the T2V model’s latent space. By blending these synthesized data with real training samples during fine-tuning, our method successfully transfers the T2V model’s conceptual knowledge to the video restoration model, thereby mitigating the distribution drift problem while preserving both texture realism and temporal coherence.
@@ 2 Related Work
Reconstruction-based Video Restoration. Early approaches focused on the architecture design and loss functions for direct HQ reconstruction from degraded inputs. CNN-based methods employed various strategies for temporal information integration, including optical flow [3, 20], deformable convolutions [32], bidirectional feature propagation [4], and optical flow-guided deformable alignment [5]. Transformer-based methods [13, 14] improved performance through attention mechanisms for long-term spatio-temporal modeling. Meanwhile, some studies [33, 42] have introduced more complex degradation simulations to improve real-world generalization. To produce richer textural details, GAN-based frameworks [34, 33] are consequently adopted that incorporate adversarial training. Despite these advances, methods relying solely on input-derived cues without strong priors still produce overly smoothed results when handling severely degraded content.
@@
Diffusion-based Video Restoration. In recent years, diffusion-based generative models [24, 22, 1, 39] have made significant progress, which introduce a new paradigm for restoration tasks. Initial explorations focused on image restoration [31, 41, 7]. Based on Stable Diffusion [24], StableSR [31] developed a trainable time-aware encoder to embed input image information into the generation process, transforming it into an image restoration model. SUPIR [41] achieved remarkable results through model scaling for image restoration. However, these approaches fundamentally lack temporal modeling capabilities, resulting in severe frame inconsistencies when directly applied to video sequences. Early solutions [46, 38] attempted to mitigate this through temporal enhancement techniques, such as incorporating trainable temporal layers [46], or implementing optical flow-based motion compensation [38], yet these post-hoc adjustments proved inadequate for ensuring robust temporal coherence. Diffusion Transformers (DiT) [21] enabled high-quality T2V generation [39] with superior temporal stability, inspiring video restoration methods. SeedVR [30] combines the shift-window attention mechanism with DiT to improve computational efficiency. STAR [36] designs a dynamic frequency loss function to improve fidelity. Concurrently, efforts to improve inference efficiency have led to the design of one-step diffusion models [29, 8]. Nevertheless, existing methods still exhibit noticeable gaps in texture realism and temporal consistency compared to native T2V models, due to distribution drift from imperfect multimodal alignment of training data. To bridge this gap, our work introduces a concept distillation training strategy that effectively preserves the texture and temproal quality of the base T2V model.
@@ 3.1 Model Architecture
ControlNet Pipeline. Given the text tokens, the visual tokens and the timestep embedding, DiT and ControlNet both perform $T$ denoising steps. DiT comprises $N$ DiT blocks, while ControlNet contains $N/7$ blocks initialized from DiT’s first $N/7$ ones. During denoising process, ControlNet’s visual tokens are integrated into DiT through $N$ proposed Dual-branch Connectors. For the $i^{th}$ connector, the fusion process of the control visual tokens is:
$\hat{f}^{i}=f^{i}+MLP(c^{\lfloor i/7\rfloor})+CA(f^{i},c^{\lfloor i/7\rfloor}),$ (1)
where $f^{i}$ denotes the visual tokens from the $i^{th}$ DiT block; $c^{\lfloor i/7\rfloor}$ represents the corresponding ControlNet block visual tokens aligned with the $i^{th}$ DiT block; $MLP(\cdot)$ and $CA(\cdot)$ are the MLP layer and cross attention module respectively; $\hat{f}^{i}$ is the fused visual tokens. After $T$ denoising steps, the visual tokens are unpatchified and fed into the VAE decoder to generate the final high-quality (HQ) outputs.
@@ 3.2 Concept Distillation Training Strategy
Training Data Collection. Effective training of DiT-based video restoration models demands extensive high-quality text-video pairs, but existing public datasets [26, 18, 37, 25] lack in both scale and diversity. To address this, we collected a large-scale video pool consisting of $3$ million videos with resolutions higher than $1024\times 1024$, frame rates higher than $24$, and durations higher than $2$ seconds. These videos cover a wide range of scenes, including portraits, natural landscapes, plants and animals, urban landscapes, etc. To ensure video quality, we further screened these videos using the no-reference video quality assessment metrics [35, 45] to remove low-quality videos. For the remaining HQ videos, we generated text descriptions using CogVLM2-Video [39], maintaining consistency with CogVideoX1.5-5B’s configuration. The final curated multimodal training dataset comprises $500K$ text-video pairs with exceptional quality and variety.
@@
Concept Distillation. Due to limitations of the VLM captioner, the constructed text-video data pairs are not perfectly aligned (see Figure 3 top row). This may lead to distribution drift during fine-tuning, degrading output video quality. While developing a more accurate VLM captioner could enhance text-video alignment, it has two drawbacks: (1) it is costly, and (2) discrepancies may persist in the T2V model’s latent space, potentially still leading to distribution drift. Instead, we address this issue by distilling the text-concept understanding capabilities of the T2V base model into the video restoration model. To this end, we employ the T2V model itself (i.e., CogVideoX1.5-5B) to perform text-guided video-to-video translation, generating training data for distillation. Specifically, given a text-video pair, we perturb the source video by adding noise with a standard deviation corresponding to $T/2$ time steps. We then apply CogVideoX1.5-5B to denoise the video over $T/2$ steps, conditioned on the text description, yielding a synthesized video with inherent alignment to text concepts in this T2V model’s latent space. As illustrated in Figure 3 (second row), the generated video largely retains the source content, but some concepts have been modified to better align with those in the text description. We randomly extract text-video pairs from the collected training dataset and employ the aforementioned process to generate $100K$ sample pairs. These generated samples are then combined with the original training dataset to facilitate fine-tuning of the control module in our DiT-based video restoration model.
@@
where $x_{0}$ is the HQ video sampled from training dataset; $x^{text}$ and $x^{lq}$ are the corresponding text description and the synthesized LQ video using the degradation model from [33]; $t$ and $\epsilon$ are the time step and noise; $x_{t}=\sqrt{\bar{\alpha_{t}}}x_{0}+\sqrt{1-\bar{\alpha_{t}}}\epsilon$ is the noised latent of $x_{0}$, and $\bar{\alpha_{t}}$ is the cumulative multiplication of the variance corresponding to time step $t$; $v_{\theta}$ denotes the networks of DiT and ControlNet (including the control feature projector and the connectors); $v$ is the optimization target, which is defined as $v=\sqrt{\bar{\alpha_{t}}}\epsilon-\sqrt{1-\bar{\alpha_{t}}}x_{0}$. During training, only the control feature projector, ControlNet, and connectors are trained, and other parameters remain frozen.
@@ 4.1 Implementation Details
Figure 5: Visual comparison results on temporal consistency. (a) displays the #5\#5 frame of the input video, and (b)-(g) present the outputs at frames 0, 55, and $10$, where “CD” denotes the proposed concept distillation. Vivid-VR demonstrates superior temporal coherence, as evident from the consistent structure of windows and doors throughout the sequence. (Zoom in on the red arrow area in each frame)
Table 2: Ablation studies of the proposed method on the UGC50 testset, where “FT” denotes “fine-tuning”, “CA” denotes “cross attention branch”, “SK” denotes “replacing MLP with a skip connection”, “QW” denotes “Qwen2.5-VL as VLM captioner”, and (h) is the setting of the proposed Vivid-VR. The best performances is marked in bold text.
@@ 4.2 Quantitative Results
To evaluate the performance of the proposed algorithm, we compare Vivid-VR against state-of-the-art approaches, including reconstruction-based methods (Real-ESRGAN [33]), generative image restoration methods (SUPIR [41]), and generative video restoration methods (UAV [46], MGLD [38], STAR [36], DOVE [8], SeedVR-7B [30], SeedVR2-7B [29]). The evaluation covers synthetic (SPMCS [27], UDM10 [40], YouHQ40 [46]) and real-world (VideoLQ [6]) benchmarks. Furthermore, we construct two testsets, containing real-world UGC videos (UGC50) and AIGC videos (AIGC50). For real-world and AIGC videos lacking of ground truth, we employed no-reference image (NIQE [17], MUSIQ [12], CLIP-IQA [28]) and video quality assessments (DOVER [35], MD-VQA [45]). For synthetic benchmarks, we supplemented these no-reference metrics with full-reference evaluations (PSNR, SSIM, and LPIPS [44]).
Table 1 presents quantitative comparisons on $6$ benchmark testsets. The proposed Vivid-VR significantly outperforms existing methods in no-reference metrics, achieving the best results in almost all metrics. At the same time, we also note its advantages in full-reference metrics appear less pronounced. We argue that this arises primarily from the inherent limitations of these metrics, which often fail to align with human perceptual preferences. For example, the LPIPS values of Figure 4(g1) and (i1) are $0.3112$ and $0.4297$, respectively, while Figure 4(i1) is more prefered by human. This phenomenon becomes particularly evident in generative restoration scenarios where severe input degradation allows multiple plausible HQ outputs, making full-reference metrics inadequate for quality assessment. This has also been mentioned in [41] and has been noticed by quality assessment studies [2, 11, 10].
@@ 4.3 Qualitative Results
In addition, Figure 5 shows visual comparisons on temporal consistency. As shown in Figure 5(g), the proposed Vivid-VR demonstrates superior coherence. For example, the structures of the windows and doors are well-consistent throughout the sequence. In contrast, SUPIR [41] shows frame-wise inconsistency as it is an image-based restoration approach (see Figure 5(c)). While STAR [36] and SeedVR-7B [30] leverage T2V frameworks, their fine-tuning-induced distribution drift compromises temporal consistency (see Figure 5(d)-(e)). Notably, Vivid-VR exhibits similar degradation when without using the proposed concept distillation strategy (see Figure 5(f)).
@@ 5 Analysis and Discussions
We have shown that the proposed Vivid-VR performs favorably against state-of-the-art methods. In this section, we perform further analysis on the key components.
@@ 5.1 Effect of the Control Feature Projector
The T2V base model is trained on HQ videos, making direct use of latent features from LQ videos detrimental to generation quality. To address this issue, we propose a control feature projector to remove degradation artifacts from the LQ video latents. To validate its effectiveness, we conducted an ablation study by disabling this module while keeping other settings unchanged. As shown in Table 2 ((a) vs (h)), the proposed control feature projector is able to improve video restoration. SUPIR [41] attempts to tackle this challenge by independently fine-tuning the VAE encoder. However, this decoupled optimization creates feature incompatibility with subsequent DiT and ControlNet, leading to suboptimal results (Table 2 (b)). While this problem can be solved through joint VAE encoder and video restoration optimization [7], this approach is expensive to train. In contrast, our proposed lightweight control feature projector can achieve similar effects at lower training cost.
Figure 6: Effect of the dual-branch connector, where “SK” denotes “replacing MLP with a skip connection”, and “CA” denotes “cross attention branch”. (Zoom-in for best view)
Figure 7: Effect of the concept distillation, where “CD” denotes “concept distillation strategy”, and “QW” denotes “Qwen2.5-VL as VLM captioner”. (Zoom-in for best view)
@@ 5.2 Effect of the Dual-Branch Connector
For the ControlNet connector, we propose a dual-branch architecture combining an MLP for feature mapping with a cross attention mechanism for dynamic feature retrieval. One may wonder whether this design helps video restoration. To answer this question, we conduct three ablation studies: 1) disabling the cross attention branch; 2) replacing the MLP branch with a skip connection; 3) adopting the ZeroSFT connector [41]. Table 2 and Figure 6 show the comparative results of the ablation experiments. When the cross attention branch is disabled, the MLP connector does not perform well and produces results lacking in detail (see Table 2(c) and Figure 6(e)). When the MLP branch is simply disabled, the video restoration model fails to converge due to its exclusive selection of DiT-like features from control inputs, resulting in output results that do not match the input content. To ensure model convergence, we replace the MLP with a skip connection. The results in Table 2(d) and Figure 6(f) show that without the MLP feature mapping, the recovered results are not well. These experiments demonstrate the necessity of our dual-branch design.
In addition, the results in Table 2(e) show that the performance of ZeroSFT connector [41] is inferior to our proposed dual-branch connector. Furthermore, the normalization operation in ZeroSFT architecture often causes residual artifacts of adjacent frames to appear in the output frames (see Figure 6(d)), while removing the normalization leads to gradient explosion during training. In contrast, our proposed connector avoids these problems.